REU 2006: SCHEDULE FOR WEEKS THREE AND FOUR

Morning talks (apprentice): Ryerson 251 Afternoon talks: Eckhart 206 MWF

Ryerson 251 TTh

Date  YSP  Abert/Babai  Abert  Babai  Farb  Fiore
Apprentice Problems  Discrete
Ryerson 251 R251/E206 R251/E206 E206/R251 E206/R251
Wed, July 5 9:30–2:30 9:30–12:00  2:45–4:00 4:30–6:00
Thu, July 6 9:30–2:30 9:30–12:00 2:45–4:15 4:30–6:00
Fri, July 7 9:30–2:30 9:30–12:00 2:45–4:00 4:30–6:00
Mon, July 10 9:30–2:30 9:30–12:00 2:45–4:15 4:30–6:00
Tue, July 11 9:30–2:30 9:30–12:00  2:45–4:00 4:30–6:00
Wed, July 12 9:30–2:30 9:30–12:00 2:45–4:15 4:30–6:00
Thu, July 13 9:30–2:30 9:30–12:00  2:45–4:00 4:30–6:00
Fri, July 14 9:30–2:30 9:30–12:00 2:45–4:15 4:30–6:00

Here is the schedule in summary form:

THIRD WEEK:

YSP: 9:30–2:30 WThF Abert/Babai (Apprentice): 9:30–12:00 WThF Abert (Problems and theorems): 2:45–4:15 Th Babai (Discrete mathematics): 4:30–6:00 Th Farb (Geometry and topology): 2:45–4:00 WF Fiore (Mathematics and music): 4:30–6:00 WF

FOURTH WEEK:

YSP: 9:30–2:30 M–F Abert/Babai (Apprentice): 9:30–12:00 M–F Abert (Problems and theorems): 2:45–4:15 MWF Babai (Discrete mathematics): 4:30–6:00 MWF Farb (Geometry and topology): 2:45–4:00 TTh Fiore (Mathematics and music): 4:30–6:00 TThu

NOTE to SESAME participants: Tutorial and work sessions should be arranged between 9:30 and 2:30. This is time free to work on problems and presentations.